OpenStudy (anonymous):

Isaac is cutting wood to build a clock. He wants the clock to have a long hand that extends from the center to the edge of the clock face. What must the approximate area of the face be if the hand he is using is 7 inches long? Use pi.

OpenStudy (anonymous):

Can you guys explain it plz?

OpenStudy (anonymous):

ya i will explain

OpenStudy (anonymous):

Area= Pi * r(squared) Area= 3.14 * (6)(squared) Area= 3.14 * 36 Area= 113.04

OpenStudy (anonymous):

If the long hand extends from the center to the edge of the clock face, then it is the circle's radius. The area of a circle is pi(r^2), where r is the radius. Therefore, the area of the clock face is pi(6)^2 = 36pi = 36(3.14) = 113.04 square inches.

OpenStudy (anonymous):

radius = 7 area = pi*r^2 = 3.14*7*7 = 3.14*49 = 153.86

OpenStudy (anonymous):

Theank mathematicalpro and facetimegeek

OpenStudy (anonymous):

im almost getting the hang of this easy stuff that i dont gett

OpenStudy (anonymous):

hes wrong

OpenStudy (anonymous):

sometimes

OpenStudy (anonymous):

i know but he tried